# MemWeaver: A Hierarchical Memory from Textual Interactive Behaviors for Personalized Generation

### v1

@@ Abstract.
The primary form of user-internet engagement is shifting from leveraging implicit feedback signals, such as browsing and clicks, to harnessing the rich explicit feedback provided by textual interactive behaviors. This shift unlocks a rich source of user textual history, presenting a profound opportunity for a deeper form of personalization. However, prevailing approaches offer only a shallow form of personalization, as they treat user history as a flat list of texts for retrieval and fail to model the rich temporal and semantic structures reflecting dynamic nature of user interests. In this work, we propose MemWeaver, a framework that weaves the user’s entire textual history into a hierarchical memory to power deeply personalized generation. The core innovation of our memory lies in its ability to capture both the temporal evolution of interests and the semantic relationships between different activities. To achieve this, MemWeaver builds two complementary memory components that both integrate temporal and semantic information, but at different levels of abstraction: behavioral memory, which captures specific user actions, and cognitive memory, which represents long-term preferences. This dual-component memory serves as a unified representation of the user, allowing large language models (LLMs) to reason over both concrete behaviors and abstracted traits. Experiments on the Language Model Personalization (LaMP) benchmark validate the efficacy of MemWeaver. Our code is available [^1] https://github.com/fishsure/MemWeaver.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
For decades, the dominant paradigm for personalization has been rooted in recommender systems that model user preferences from implicit feedback signals, such as clicks and purchases (Hu et al., 2008; Oard et al., 1998; Cheng et al., 2021). In this paradigm, users are often represented as vectors in a latent space, derived from their interactions with a catalog of item IDs(Zhang et al., 2025b). However, a fundamental shift is underway, with the primary mode of user engagement evolving from passive content consumption guided by implicit behaviors to the active co-creation of information through explicit textual conversations (Gao et al., 2021; Sun et al., 2019; Zhao et al., 2024). As users increasingly engage in these direct interactive behaviors, the volume and depth of their textual history grow, creating a new foundation for deep user modeling and personalization (Geng et al., 2022; Qin et al., 2025).
@@
Despite this ongoing paradigm shift, much of the existing user modeling approach remains reliant on implicit signals (Wang et al., 2019; He et al., 2020; Cheng et al., 2024b). Classical approaches, such as collaborative filtering and matrix factorization, construct user profiles by learning embeddings from a history of interactions with atomic item identifiers (IDs) (Kang and McAuley, 2018; Cheng et al., 2024a). While effective for capturing broad preference patterns from sparse data, these ID-based models have an inherent limitation: they are not designed to interpret the rich semantic content of language. The user’s intent is learned indirectly through collaborative patterns, rather than being understood directly from their own words, leaving the nuanced details of their goals largely unexplored.
@@
The emergence of this rich textual history creates a profound opportunity for a deeper form of personalization (Liu et al., 2025; Cheng et al., 2024c). This history offers a direct, high-fidelity narrative of a user’s evolving interests, nuanced intentions, and specific preferences, far exceeding the expressive power of item IDs. However, harnessing this information is not trivial. The true value of this history lies not in its sheer volume, but in its intricate nature, embodying two critical dimensions: the temporal evolution of a user’s interests and the latent semantic connections that link disparate topics (Shi et al., 2025; Salemi et al., 2024; Cheng et al., 2022). The central challenge, therefore, is to develop a framework capable of harnessing both the temporal evolution and the semantic connections within this history to model the user (Cheng et al., 2025).
@@
To address this challenge, we introduce MemWeaver, a framework designed to weave a hierarchical memory from user interaction history. To capture both the temporal evolution of interests and the semantic relationships between user behaviors, MemWeaver constructs a dual-component memory comprising a behavioral memory that captures concrete actions and a cognitive memory that abstracts higher-level intent. To form the behavioral memory, we first represent the user’s history as a network where each behavior is a node, interconnected by two types of edges: temporal edges link consecutive behaviors, while semantic edges connect those with high thematic relevance (Edge et al., 2024; Li et al., 2025b). The memory is then extracted from this network via a context-aware random walk (Pearson, 1905) that simulates the associative recall process of human memory, traversing the established edges. Concurrently, the cognitive memory is abstracted through a process that segments the history into distinct phases, summarizes the interests within each, and integrates these summaries into a global profile of user preferences.
@@ 2.1. User Modeling for Personalization
The field of user modeling for personalization is rooted in traditional recommender systems, where the existing methods involve modeling preferences from sparse user-item interaction matrices using techniques such as collaborative filtering (CF) and matrix factorization (Koren et al., 2009; Rendle et al., 2009). While effective, these models operate on atomic item identifiers (IDs) and remain oblivious to the underlying semantic content, learning preferences only indirectly. The pursuit of capturing dynamic user interests has led to an evolution toward sequential modeling with RNNs and Transformers (Hidasi et al., 2015; Kang and McAuley, 2018; Sun et al., 2019), and later to graph-based approaches using graph neural networks (GNNs) to capture higher-order relationships (Wang et al., 2019; He et al., 2020). Despite increasing improvement in modeling temporal patterns and relational structures, these methods share a fundamental limitation: they are tailored for sparse, ID-level signals and are ill-equipped to interpret the rich semantics of textual interactions. The advent of LLMs introduces a new frontier, shifting the focus to directly understanding user narratives from textual history via retrieval-augmented generation (RAG) (Lewis et al., 2020; Zhu et al., 2025). However, the efficacy of current RAG-based personalization is hampered by simplistic, single-dimensional retrieval heuristics, such as recency-based or similarity-based approaches (Karpukhin et al., 2020; Reimers and Gurevych, 2019), and their treatment of retrieved context as a flat, unstructured list of text chunks. This approach fails to fully exploit the LLM’s reasoning capabilities, as it strips away the intricate relationships within a user’s history. A critical gap therefore lies in the absence of a framework capable of constructing and leveraging a hierarchical memory from textual interaction histories.
@@ 2.2. Memory-Augmented LLMs
Recent studies have increasingly investigated memory mechanisms in LLMs, aiming to enhance their capabilities in long-context understanding, multi-turn reasoning, and user-centric personalization (Zhang et al., 2024a; Chhikara et al., 2025; Zhong et al., 2024). To support long-context modeling, prior work extends the context window or compresses historical information efficiently. MemoRAG (Qian et al., 2025) and Memory3 (Yang et al., 2024) apply key–value compression to fit more content within fixed memory budgets, while Mem-Agent (Yu et al., 2025) and Memory-R1 (Yan et al., 2025) use reinforcement learning to guide summarization policies that preserve salient information across long documents. When it comes to complex reasoning and interaction, memory modules have been used to support multi-hop inference and agentic decision-making. HippoRAG (Jimenez Gutierrez et al., 2024) leverages knowledge graphs to emulate associative reasoning across linked facts. Memento and Agent-KB (Tang et al., 2025) incorporate memory to reuse past successes for in-context decisions. Other systems, such as Mirix (Wang and Chen, 2025) and Mem-OS (Li et al., 2025a), build complex memory architectures integrating multiple types of memory for more general-purpose agent behaviors. For personalized generation, recent studies incorporate memory to model user-specific knowledge and preferences. Prime (Zhang et al., 2025a) introduces a dual-memory structure with chain-of-thought reasoning to track evolving preferences, while MiLP (Zhang et al., 2024b) leverages multiple LoRA modules as parametric memory units, selected via Bayesian optimization to adaptively capture user-specific signals. Despite these advances, the structural modeling of user interaction histories remains underexplored. In particular, existing methods often fail to jointly model the semantic and temporal relationships inherent in user histories. To address this limitation, we propose MemWeaver, a framework that constructs a hierarchical memory to effectively capture both the semantic and temporal information within user history.
@@ 3.3. Memory Construction and Updating
We first detail the construction process for behavioral memory and cognitive memory and then describe the mechanisms for their incremental update strategies.
@@ 3.3.2. Cognitive Memory Abstraction
3.3.3. Incremental Memory Updating
To ensure our framework can efficiently handle continuously growing user histories without costly full reconstructions, we employ lightweight incremental update mechanisms for both memory components. For the behavioral memory, when a new batch of behaviors $\mathcal{H}_{\text{new}}$ arrives, it is encoded and integrated without reprocessing the existing graph. Temporal continuity is maintained by linking the last existing node to the first new one, and new semantic edges are established by clustering the new behaviors independently. Similarly, for the cognitive memory, the new batch is processed independently by partitioning it into coherent segments and generating new local summaries. The global cognitive memory is then re-synthesized by prompting the LLM with the full collection of local summaries, combining the previously generated ones with the new ones. This approach requires only a single, high-level integration step to update the user’s overall narrative, making the process computationally tractable.
@@ 4.1. Experimental Setup
Our experimental setup is as follows: we first introduce the LaMP benchmark dataset, then specify the evaluation metrics for each task, and finally describe the baseline methods used for comparison.
@@ 4.1.1. Dataset
We adopt the LaMP benchmark (Salemi et al., 2023), which covers seven personalized tasks, including three classification tasks—LaMP-1 (Citation Identification, binary), LaMP-2 (Movie Tagging, multi-class with 15 categories), and LaMP-3 (Product Rating, ordinal from 1–5 stars)—and four text generation tasks: LaMP-4 (News Headline Generation), LaMP-5 (Scholarly Title Generation), LaMP-6 (Email Subject Generation), and LaMP-7 (Tweet Paraphrasing). Since LaMP-6 is not publicly available, we exclude it from our experiments. Following the official protocol, we adopt a time-based split, where each user’s interactions are partitioned into training, validation, and test sets by timestamp. This setting simulates a realistic scenario and allows us to study the effect of recency.
@@
Main
Component
w/o Cognitive Memory 0.6417 0.3209 0.4525 0.0301 0.2988 0.3871 0.1651 0.1477 0.4630 0.4076 0.4543 0.3817
@@
w/ Behavioral
Memory
w/o Temporal Edges 0.6233 0.3117 0.4417 0.0291 0.3016 0.4084 0.1636 0.1461 0.4621 0.4032 0.4488 0.3771
@@
w/ Cognitive
Memory
w/o Clustering 0.6431 0.3216 0.4546 0.0302 0.2892 0.3710 0.1665 0.1487 0.4625 0.4052 0.4620 0.3893
@@ 4.1.4. Implementation Details
For all experiments, we use Qwen3-8B (Yang et al., 2025) and Llama-3.1-8B-Instruct (Dubey et al., 2024) as the backbone LLMs to evaluate our framework and all compared baselines. Unless otherwise specified, results are reported using Qwen3-8B. Further implementation details are provided in Appendix A.
@@ 4.3.2. Analysis of Behavioral Memory Components
We performed a fine-grained ablation to dissect the contributions of the components within the behavioral memory. The primary observation is that the complete model consistently outperforms all its ablated variants, confirming that temporal edges, semantic edges, and the edge weighting mechanism are all integral to its performance. When comparing the two types of connections, we find that the removal of semantic edges causes a more significant performance drop than the removal of temporal edges. This suggests that the semantic content and relationships between user actions carry more weight in constructing an effective user profile than their purely sequential ordering. Finally, the study reveals that the edge weighting mechanism is the most critical component. The variant without edge weighting suffers the most severe performance degradation, demonstrating that simply constructing a graph of behaviors is insufficient. Instead, the ability to quantify the relevance of these relationships through fine-grained weighting is critically important.
@@ 4.3.3. Analysis of Cognitive Memory Strategies
To investigate the contributions of the components within the cognitive memory, we performed a fine-grained ablation study. First, our findings confirm the importance of our hierarchical methodology. Removing the initial segmentation and clustering component leads to a significant performance degradation. This underscores the benefits of our approach: by partitioning the user’s history into temporally coherent phases and summarizing them individually, our model can capture the distinct characteristics of the user’s interests at different time periods. This staged process allows the final cognitive memory not only to achieve high-level semantic abstraction but also to effectively model the evolution of the user’s interests over time. Furthermore, it avoids the pitfalls of summarizing the entire user history at once, which could lead to an excessively long context that degrades the quality of the resulting summary. The results also confirm the efficacy of the final aggregation step, as the model variant without a global summary shows a consistent drop in performance, demonstrating that effectively distilling these temporal summaries into a single, high-level profile is critical for capturing the user’s core traits.
@@ 4.4.1. Analysis of Edge Traversal Proportions
To further understand the interplay between these edge types during inference, we analyzed the proportion of edges traversed by the random walk on different datasets, as illustrated in Figure 3. A key finding is that the traversal preference varies significantly based on the inherent characteristics of the dataset. For instance, in a task highly sensitive to chronological context, such as news headline generation, temporal edges constitute the vast majority of the traversed paths. Conversely, for citation identification, where recognizing related work is essential, semantic edges are traversed with much greater frequency. This demonstrates our model’s ability to dynamically adapt its focus according to the nature of the task.
Crucially, we observe from the figure that temporal edges account for a majority of traversals across all datasets. This dominance of temporal edges underscores the fundamental importance of the temporal dimension in modeling user behavior and presents a strong counterargument to many existing methods that neglect chronological factors. Interestingly, while temporal edges are traversed more frequently, the ablation study indicates that removing semantic edges incurs a greater performance penalty. This suggests that temporal edges represent the high-frequency paths of user behavior, whereas semantic edges, though traversed less often, act as essential links connecting conceptually related but chronologically distant activities. Consequently, removing these vital semantic links compromises the graph’s integrity more severely, thus explaining the greater performance penalty.

### v2

@@ Abstract.
The primary form of user-internet engagement is shifting from leveraging implicit feedback signals, such as browsing and clicks, to harnessing the rich explicit feedback provided by textual interactive behaviors. This shift unlocks a rich source of user textual history, presenting a profound opportunity for a deeper form of personalization. However, prevailing approaches offer only a shallow form of personalization, as they treat user history as a flat list of texts for retrieval and fail to model the rich temporal and semantic structures reflecting dynamic nature of user interests. In this work, we propose MemWeaver, a framework that weaves the user’s entire textual history into a hierarchical memory to power deeply personalized generation. The core innovation of our memory lies in its ability to capture both the temporal evolution of interests and the semantic relationships between different activities. To achieve this, MemWeaver builds two complementary memory components that both integrate temporal and semantic information, but at different levels of abstraction: behavioral memory, which captures specific user actions, and cognitive memory, which represents long-term preferences. This dual-component memory serves as a comprehensive representation of the user, allowing large language models (LLMs) to reason over both concrete behaviors and abstracted cognitive traits. This leads to content generation that is deeply aligned with their latent preferences. Experiments on the six datasets of the Language Model Personalization (LaMP) benchmark validate the efficacy of MemWeaver. Our code is available [^1] https://github.com/fishsure/MemWeaver.
@@
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emirates.booktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesisbn: 979-8-4007-2307-0/2026/04doi: 10.1145/3774904.3792732ccs: Computing methodologies Natural language generationccs: Computing methodologies Information extractionccs: Computing methodologies Discourse, dialogue and pragmatics
@@ 1. Introduction
For decades, the dominant paradigm for personalization has been rooted in recommender systems that model user preferences from implicit feedback signals, such as clicks and purchases (Hu et al., 2008; Oard et al., 1998; Cheng et al., 2021; Wang et al., 2025b, c). In this paradigm, users are often represented as vectors in a latent space, derived from their interactions with a catalog of item IDs(Zhang et al., 2025b; Liu et al., 2022). However, a fundamental shift is underway, with the primary mode of user engagement evolving from passive content consumption guided by implicit behaviors to the active co-creation of information through explicit textual conversations (Gao et al., 2021; Sun et al., 2019; Zhao et al., 2024). As users increasingly engage in these direct interactive behaviors, the volume and depth of their textual history grow, creating a new foundation for deep user modeling and personalization (Geng et al., 2022; Qin et al., 2025).
@@
Despite this ongoing paradigm shift, much of the existing user modeling approach remains reliant on implicit signals (Wang et al., 2019; He et al., 2020). Classical approaches, such as collaborative filtering and matrix factorization, construct user profiles by learning embeddings from a history of interactions with atomic item identifiers (IDs) (Kang and McAuley, 2018). While effective for capturing broad preference patterns from sparse data, these ID-based models have an inherent limitation: they are not designed to interpret the rich semantic content of language. The user’s intent is learned indirectly through collaborative patterns, rather than being understood directly from their own words, leaving the nuanced details of their goals largely unexplored.
@@
The emergence of this rich textual history creates a profound opportunity for a deeper form of personalization (Liu et al., 2025a; Cheng et al., 2026). This history offers a direct, high-fidelity narrative of a user’s evolving interests, nuanced intentions, and specific preferences, far exceeding the expressive power of item IDs. However, harnessing this information is not trivial. The true value of this history lies not in its sheer volume, but in its intricate nature, embodying two critical dimensions: the temporal evolution of a user’s interests and the latent semantic connections that link disparate topics (Shi et al., 2025; Salemi et al., 2024). The central challenge, therefore, is to develop a framework capable of harnessing both the temporal evolution and the semantic connections within this history to model the user.
@@
To address this challenge, we introduce MemWeaver, a framework designed to weave a hierarchical memory from user interaction history. To capture both the temporal evolution of interests and the semantic relationships between user behaviors, MemWeaver constructs a dual-component memory comprising a behavioral memory that captures concrete actions and a cognitive memory that abstracts higher-level intent. To form the behavioral memory, we first represent the user’s history as a network where each behavior is a node, interconnected by two types of edges: temporal edges link consecutive behaviors, while semantic edges connect those with high thematic relevance (Edge et al., 2024; Li et al., 2025a). The memory is then extracted from this network via a context-aware random walk (Pearson, 1905) that simulates the associative recall process of human memory, traversing the established edges. Concurrently, the cognitive memory is abstracted through a process that segments the history into distinct phases, summarizes the interests within each, and integrates these summaries into a global profile of user preferences.
@@ 2.1. User Modeling for Personalization
Traditional personalization methods evolved from ID-based collaborative filtering and matrix factorization (Koren et al., 2009; Rendle et al., 2009; Liu et al., 2025b, 2024) to modeling with RNNs and Transformers (Kang and McAuley, 2018; Sun et al., 2019) and graph-based models (Wang et al., 2019; He et al., 2020). Despite increasing improvement in modeling temporal patterns and relational structures, these methods share a fundamental limitation: they are tailored for sparse, ID-level signals and are ill-equipped to interpret the rich semantics of textual interactions (Zhang et al., 2025a, 2023). The advent of LLMs introduces a new frontier, shifting the focus to directly understanding user narratives from textual history via retrieval-augmented generation (RAG) (Lewis et al., 2020; Zhu et al., 2025; Yu et al., 2025b). However, the efficacy of current RAG-based personalization is hampered by simplistic, single-dimensional retrieval heuristics, such as recency-based or similarity-based approaches (Cheng et al., 2025; Karpukhin et al., 2020; Reimers and Gurevych, 2019), and their treatment of retrieved context as a flat, unstructured list of text chunks. Discarding relational context fails to fully exploit LLM reasoning, revealing a critical gap: the lack of a framework to derive hierarchical memory from textual interaction histories.
@@ 2.2. Memory-Augmented LLMs
Recent studies have investigated memory mechanisms to boost LLMs in long-context understanding, multi-turn reasoning, and personalization (Zhang et al., 2024b; Chhikara et al., 2025; Zhong et al., 2024). To facilitate long-context modeling, prior works either extend context windows or compress information: MemoRAG (Qian et al., 2025) and Memory3 (Yang et al., 2024) apply KV-compression, while Mem-Agent (Yu et al., 2025a) and Memory-R1 (Yan et al., 2025) utilize reinforcement learning to refine summarization policies that preserve salient information. For complex reasoning, memory modules support multi-hop and agentic decision-making; for instance, HippoRAG (Jimenez Gutierrez et al., 2024) emulates associative reasoning via knowledge graphs, while Memento and Agent-KB (Tang et al., 2025; Wang et al., 2025a) store past experiences for in-context reuse. Further, Mirix (Wang and Chen, 2025) and Mem-OS (Li et al., 2025b) integrate heterogeneous memory types to support general-purpose agent behaviors. For personalization, Prime (Zhang et al., 2025c) employs a dual-memory structure with chain-of-thought, whereas MiLP (Zhang et al., 2024a) uses Bayesian-optimized LoRA modules as parametric memory to capture user-specific signals. Despite these advances, existing methods often fail to jointly model the intertwined semantic and temporal structures in user histories. To address this limitation, we propose MemWeaver, a framework that constructs a hierarchical memory to capture both the semantic and temporal information within user history.
@@ 3.3.2. Cognitive Memory Abstraction
3.3.3. Incremental Memory Updating
To ensure our framework can efficiently handle continuously growing user histories without costly full reconstructions, we employ lightweight incremental update mechanisms for both memory components. For the behavioral memory, when a new batch of behaviors $\mathcal{H}_{\text{new}}$ arrives, it is encoded and integrated without reprocessing the existing graph. Temporal continuity is maintained by linking the last existing node to the first new one, and new semantic edges are established by clustering the new behaviors independently. Similarly, for the cognitive memory, the new batch is processed independently by partitioning it into coherent segments and generating new local summaries. The global cognitive memory is then re-synthesized by prompting the LLM with the full collection of local summaries, combining the previously generated ones with the new ones. This approach requires only a single, high-level integration step to update the user’s overall narrative, making the process computationally tractable. Operating on condensed summaries bounds token costs as history grows, enabling real-time adaptation to evolving interests without the latency of full-context re-computation.
@@ 4.1.1. Dataset
We adopt the LaMP benchmark (Salemi et al., 2023), which covers seven personalized tasks, including three classification tasks—LaMP-1 (Citation Identification, binary), LaMP-2 (Movie Tagging, multi-class with 15 categories), and LaMP-3 (Product Rating, ordinal from 1–5 stars)—and four text generation tasks: LaMP-4 (News Headline Generation), LaMP-5 (Scholarly Title Generation), LaMP-6 (Email Subject Generation), and LaMP-7 (Tweet Paraphrasing). Since LaMP-6 is not publicly available, we exclude it from our experiments. Following the official protocol, we adopt a time-based split, where each user’s interactions are partitioned into training, validation, and test sets by timestamp. This setting simulates a realistic scenario and allows us to study the effect of recency. To demonstrate the generalizability of our approach, we conduct additional experiments on other datasets which are detailed in Appendix C.
@@
Main Component w/o Cognitive Memory 0.6417 0.3209 0.4525 0.0301 0.2988 0.3871 0.1651 0.1477 0.4630 0.4076 0.4543 0.3817
@@
w/ Behavioral Memory w/o Temporal Edges 0.6233 0.3117 0.4417 0.0291 0.3016 0.4084 0.1636 0.1461 0.4621 0.4032 0.4488 0.3771
@@
w/ Cognitive Memory w/o Clustering 0.6431 0.3216 0.4546 0.0302 0.2892 0.3710 0.1665 0.1487 0.4625 0.4052 0.4620 0.3893
@@ 4.1.4. Implementation Details
For all experiments, we use Qwen3-8B (Yang et al., 2025) and Llama-3.1-8B-Instruct (Dubey et al., 2024) as the backbone LLMs to evaluate our framework and all compared baselines. We utilize BGE-M3 (Chen et al., 2024) as the retrieval model. Unless otherwise specified, results are reported using Qwen3-8B. Further implementation details are provided in Appendix A.
@@ 4.3.2. Analysis of Behavioral Memory Components
We performed a fine-grained ablation to dissect the contributions of the components within the behavioral memory. The complete model consistently outperforms all its ablated variants, confirming that temporal edges, semantic edges, and the edge weighting mechanism are all integral to its performance. When comparing the two types of connections, we find that the removal of semantic edges causes a more significant performance drop than the removal of temporal edges. This suggests that the semantic content and relationships between user actions carry more weight in constructing an effective user profile than their purely sequential ordering. Finally, the study reveals that the edge weighting mechanism is the most critical component. The variant without edge weighting suffers the most severe performance degradation, demonstrating that simply constructing a graph of behaviors is insufficient. Instead, the ability to quantify these relationships through fine-grained weighting is important.
@@ 4.3.3. Analysis of Cognitive Memory Strategies
To investigate the contributions of the components within the cognitive memory, we performed a fine-grained ablation study. Findings underscore the importance of our hierarchical methodology. Removing the initial segmentation and clustering component leads to a significant performance degradation. This underscores the benefits of our approach: by partitioning the user’s history into temporally coherent phases and summarizing them individually, our model can capture the distinct characteristics of the user’s interests at different time periods. This staged process allows the final cognitive memory not only to achieve high-level semantic abstraction but also to effectively model the evolution of the user’s interests over time. Furthermore, it avoids the pitfalls of summarizing the entire user history at once, which could lead to an excessively long context that degrades the quality of the resulting summary. Results also confirm the efficacy of the final aggregation step, as the model variant without a global summary shows a consistent drop in performance, demonstrating that effectively distilling these temporal summaries into a single, high-level profile is critical for capturing the user’s core traits.
@@ 4.4.1. Analysis of Edge Traversal Proportions
Figure 3 illustrates the proportion of edges traversed during inference, revealing that the model dynamically adapts to task characteristics. In chronological tasks like news generation, temporal edges constitute the vast majority of paths, while semantic edges gain significant prominence in association-driven domains like citation identification. Crucially, temporal edges account for the majority of traversals across all datasets, validating the fundamental importance of the temporal dimension often neglected by existing methods. Interestingly, while temporal edges are traversed more frequently, removing semantic edges incurs a larger performance penalty. This disparity suggests that temporal edges form the high-frequency backbone of user behavior, whereas semantic edges act as essential bridges connecting conceptually related but temporally distant activities. Consequently, removing these vital semantic links compromises the graph’s structural integrity more severely.